There are three exams in a marking period. A student received 75 and 80 on the first two exams. What grade must the student earn on the last exam to get an average of no less than 80% for the marking period.
Answer by jorel1380(3719)

3x80=240
75+80+n=240
n=240-155=85
The student needs no less than an 85..